OpenStudy (anonymous):

How do you find the P-Value? Ho: P=0.6 versus H1: P>0.6 N=250, X=170, A=0.1

OpenStudy (zarkon):

\[\Huge z=\frac{\hat{p}-p_0}{\sqrt\frac{p_0(1-p_0)}{n}}\]

OpenStudy (zarkon):

Then compute \[P(Z>z)\] Where \(Z\sim N(0,1)\)

OpenStudy (anonymous):

Ok Thank you,I'll use that formula and work it out.